# Baking Gaussian Splatting into Diffusion Denoiser for Fast and Scalable Single-stage Image-to-3D Generation

### v2

@@ 4.2 Ablation Study
Break-down Ablation. To study the effect of each component towards higher performance, we adopt the denoiser without timestep control as the baseline to conduct a break-down ablation. We train it on the object-level datasets with a single-view input and the same amount of supervised views as DiffusionGS. Results on GSO [13] are reported in Tab. 2. The baseline yields poor results of 17.63 dB in PSNR and 118.31 in FID. When applying our diffusion framework, point cloud distribution loss $\mathcal{L}_{pd}$ in Eq.(12), scene-object mixed training without the reference-point plücker coordinate (RPPC), and RPPC, the model gains by 2.94, 0.37, 0.79, 0.34 dB in PSNR and drops by 70.45, 19.45, 10.62, 6.27 in FID. Besides, the improvement of RPPC on RealEstate10K test set is 0.28 dB in PSNR and 7.09 in FID. These results suggest the effectiveness of our methods.
Analysis of Mixed Training. We conduct a visual analysis of our scene-object mixed training in Fig. 8 (a). For fair comparison, models are trained with the same number of iterations whether with or without mixed training. The upper row shows the effect on object-level generation. After using the mixed training, the textures of the cup become clearer and more realistic, and the artifacts on the back are reduced. The lower row depicts the effect on scene-level generation. When applying our mixed training, DiffusionGS can better capture the 3D geometry and generate more structural contents such as the window and stove of the kitchen.
Analysis of Generation Diversity. We evaluate the generation diversity of DiffusionGS by changing the random seed with the same front prompt view “ikun” as shown in the second row of Fig. 1. Different generated samples are shown in Fig. 8 (b). DiffusionGS can generate different shapes and textures for 3D assets, such as the back of the “ikun” doll.
@@ 5 Conclusion
In this paper, we propose a novel 3DGS-based diffusion model, DiffusionGS, for single-stage object and scene generation from a single view. Our DiffusionGS directly outputs 3D Gaussian point clouds at each timestep to enforce view consistency and only requires 2D renderings for supervision. In addition, we develop a scene-object mixed training strategy with a new camera conditioning method RPPC to learn a general prior capturing better 3D geometry and texture representations. Experiments demonstrate that our DiffusionGS significantly outperforms the SOTA 3D diffusion model by 2.2/2.91 dB in PSNR and 23.25/75.68 in FID on object/scene generation while enjoying over 5$\times$ speed ($\sim$6s) on single A100 GPU. The user study and text-to-3D application also reveal the practical values of our method.

### v3

@@ 4.2 Ablation Study
Break-down Ablation. To study the effect of each component towards higher performance, we adopt the denoiser without timestep control as the baseline to conduct a break-down ablation. We train it on the object-level datasets with a single-view input and the same amount of supervised views as DiffusionGS. Results on GSO [13] are reported in Tab. 2. The baseline yields poor results of 17.63 dB in PSNR and 118.31 in FID. When applying our diffusion framework, loss $\mathcal{L}_{pd}$ in Eq.(12), scene-object mixed training without RPPC, and RPPC, the model gains by 2.94, 0.37, 0.79, 0.34 dB in PSNR and drops by 70.45, 19.45, 10.62, 6.27 in FID.
Analysis of Mixed Training. We conduct an analysis of our scene-object mixed training in Fig. 9 (a). For fair comparison, models are trained with the same iterations whether with or without mixed training. (i) The left part shows the effect on object generation. After using the mixed training, the textures of the cup become clearer and more realistic, and the artifacts on the back are reduced. (ii) The right part depicts the effect on scene reconstruction. When applying our mixed training, DiffusionGS can better capture the 3D geometry and reconstruct the house with less artifacts. The improvement of our mixed training for scene reconstruction on Realestate10K is 0.61 dB in PSNR and 10.53 in FID.
Analysis of RPPC. (i) Using RPPC leads to a quantitative improvement of 0.28 dB in PSNR and 7.09 in FID for scene reconstruction on Realestate10K. (ii) Fig. 8 (c) and (d) analyze the effect of RPPC. Using RPPC can render more accurate scene structures (window, floor, etc.) and depth.
Analysis of Generation Diversity. We change the random seed with the same prompt view in Fig. 9 (b). Our method can generate different shapes and textures for 3D assets.
@@ 5 Conclusion
In this paper, we propose a novel framework, DiffusionGS, for 3D object generation and scene reconstruction from a single view. Our DiffusionGS directly outputs 3D Gaussian point clouds at each timestep to enforce view consistency and only requires 2D renderings for supervision. In addition, we develop a scene-object mixed training strategy with a new camera conditioning method RPPC to learn a general prior capturing better 3D geometry and texture representations. Experiments show that our method outperforms SOTA methods while enjoying a faster speed of 6 seconds.